Susan has 600 feet of fencing to use in making a yard for her new puppy. The yard she has planned has a length that is 75 feet less than twice the width. What are the dimensions of the yard?
Answer by alicealc(293)

Length = L
Width = W

length is 75 feet less than twice the width:
L = 2W - 75
perimeter = 600 feet
perimeter = 2 * (length + width)
600 = 2 * (2W - 75 + W)
600/2 = 2W - 75 + W
300 = 3W - 75
300 + 75 = 3W
375 = 3W
W = 375/3 = 125
L = 2W - 75 = 2 * 125 - 75 = 250 - 75 = 175

so the dimension of the yard is:
length = 175 feet
width = 125 width